# Bayesian inversion with Student’s $t$ priors based on Gaussian scale mixtures

### v1

@@ 2.3 Gaussian scale mixture representation
where the density $p_{\scriptscriptstyle X|Z}(x\;|\;z)$ is conditionally Gaussian given a realization $z$. Essentially, the GSM density involves a marginalization of the mixing random variable $Z$. In some cases this marginalization is analytical. For example, several continuous distributions can be constructed using a GSM representation: the logistic and Laplace distributions Andrews and Mallows (1974), generalized Gaussian distributions West (1987), as well as, the Student’s $t$ and $\alpha$-stable families of distributions (Feller, 1971, p.173).
When GSMs are used as priors for Bayesian inference, they induce a hierarchical structure that assumes Gaussianity at the first layer, and the mixing random variable becomes a hyperparameter at the second layer (see, e.g., Choy and Smith (1997); Wainwright and Simoncelli (1999)). This representation is ideal in iterative algorithms such as Markov chain Monte Carlo (MCMC), since upon the introduction of the auxiliary mixing variable, the target random variable becomes conditionally Gaussian. Standard and efficient techniques are available to deal with Gaussian distributions compared to those required for more involved sub-Gaussian distributions.
@@ 3 Student’s $t$ MRF prior based on Gaussian scale mixture representation
We introduce a prior that integrates the MRF structure with the parametric family of Student’s $t$ distributions. This distribution is characterized by tail-heaviness, which is modulated through the degrees of freedom parameter. Subsequently, we outline a method to explore the posterior distribution under this prior, with a particular focus on the modeling of the degrees of freedom parameter.
@@ 3.1 Formulation of the prior
We investigate a parametric family of MRF priors designed to capture both sharp and smooth features. This family utilizes the Gosset distribution, commonly known as the Student’s $t$ distribution or simply the $t$ distribution (we use this names interchangeably throughout the presentation). We employ this general family as a prior for each component in the difference random vector
@@
where $\mathcal{T}_{\nu}\left(\mu,\tau^{2}\right)$ denotes the Student’s $t$ distribution, and it has density function (Feller, 1971, p.49)
@@
Notably, in the limit $\nu\to\infty$, the $t$ distribution approaches a Gaussian distribution. Conversely, when $\nu=1$, the distribution corresponds to a Cauchy distribution. It is worth noting that in practical applications, the degrees of freedom parameter is typically constrained to values greater than one or two, ensuring a well-defined mean or variance for the prior distribution, respectively. Consequently, to emphasize edge preservation, attention is directed towards distributions with lower degrees of freedom; in the opposite case, when aiming to represent smooth features, distributions with higher degrees of freedom become necessary. To illustrate the relationship between the degrees of freedom and the tail-heaviness of a $t$ distribution, we visually depict in Figure 1 the logarithm of the density function of the univariate standard $t$ distribution $\mathcal{T}_{\nu}\left(0,1\right)$ with increasing degrees of freedom.
Figure 1: Logarithm of the probability density of the univariate standard Student’s $t$ distribution $\mathcal{T}_{\nu}\left(0,1\right)$ for different values of the degrees of freedom parameter $\nu$. As $\nu=\infty$, Student’s $t$ $\mathcal{T}_{\infty}\left(0,1\right)$ is equivalent to the univariate standard Gaussian distribution $\mathcal{N}(0,1)$.
Sampling from posterior distributions in high dimensions under the Student’s $t$ MRF prior is cumbersome. To improve the efficiency of the prior implementation one can utilize the GSM representation of the Student’s $t$ distribution. In this case, the mixing distribution is inverse-gamma Andrews and Mallows (1974). By incorporating this property, we can transform the original prior model into a hierarchical one, where the target difference random vector follows (conditionally) a more tractable Gaussian distribution. Therefore, we can write the GSM prior density for the $i$-th component of difference random vector as follows
@@
where ${w}_{i}^{2}$ is the mixing random variable and $\text{IG}(a,b)$ denotes the inverse-gamma distribution with shape and scale parameters $a,b$, such that the density is defined as
$p(w_{i}^{2})=\frac{b^{a}}{\Gamma(a)}(w_{i}^{2})^{-a-1}\exp\left(-\frac{b}{w_{i}^{2}}\right).$ (11)
Note that each mixing variable $w_{i}^{2}$ appears to control the variance locally for each difference component $u_{i}$. Analogously, the scale-squared parameter $\tau^{2}$ controls the prior variance globally across the domain. As this parameter becomes important for the prior specification, we assume it is uncertain and endow it with another inverse-gamma hyperprior. In this case, we set the shape parameter as $a=1$ and the scale parameter as $b=10^{-4}$ (see, e.g., Higdon (2007)).
By utilizing GSM in eq. 10b and the linear relation described in eq. 5, we can establish a conditionally Gaussian MRF prior for the random vector $\bm{X}$. In Gaussian MRFs, the precision matrix plays a crucial role, as its sparsity structure determines the neighborhood system that explains conditional relationships among the elements of $\bm{X}$. In edge-preserving inverse problems, the precision matrix is constructed based on the differences between the elements of the parameter vector. Specifically, it depends on the matrix $\bm{L}$ (for detailed information, refer to (Bardsley, 2019, p.68)). Consequently, we can express the prior precision matrix as
@@
As a result, by using the GSM of the $t$ distribution on the random vector of differences $\bm{U}$, we establish a hierarchical prior for $\bm{X}$ as follows
$p(\bm{x},\tau^{2},\bm{w}^{2})=p(\bm{x}\;|\;\tau^{2},\bm{w}^{2})p(\bm{w}^{2}\;|\;\nu)p(\nu)p(\tau^{2}),$ (13a)
@@
$\tau^{2}\sim\text{IG}\left(a,b\right),\quad\bm{w}^{2}\;|\;\nu\sim\prod_{i=1}^{d}\text{IG}\left(\nicefrac{{\nu}}{{2}},\nicefrac{{\nu}}{{2}}\right),\quad\nu\sim p(\nu),$ (13c)
@@
The global-local scheme associated with the hierarchical prior eq. 13a shares similarities with the horseshoe prior Uribe et al. (2023), however, the hyperpriors for the local and global scale parameters are defined as half-Cauchy distributions for which the GSM representation has a different form. Moreover, the hierarchical prior eq. 13a is also related to the models discussed in Calvetti et al. (2020); Law and Zankin (2022), which employ a generalized inverse-gamma distribution as hyperprior.
@@ 3.2 Modeling the degrees of freedom
The estimation of the degrees of freedom parameter $\nu$ of the Student’s $t$ distribution has been actively studied Geweke (1993); Fernandez and Steel (1998); Fonseca et al. (2008). The choice of the prior distribution is known to be challenging as this parameter is typically poorly identified from the data Villa and Walker (2013).
We consider two popular prior choices Juárez and Steel (2010); Lee (2022):
@@
    $\text{Ga}(\nu;\,a=2,b=0.1,\mu=0)$, the gamma distribution with the shape parameter $a\in\mathbb{R}_{>0}$, rate parameter $b\in\mathbb{R}_{>0}$ and location parameter $\mu\in\mathbb{R}_{>0}$. The gamma probability density is defined as
@@
    $\text{Ga}(\nu;\,a=2,b=0.1,\mu=1)$, a thresholded Gamma distribution which is equivalent to the prior choice $1$, but now excluding values $\nu\leq 1$. We remark that in this case $\mu\neq 0$, and we further employ notation $\mathrm{Ga}_{\nu>\mu}(\nu;\,a,b)$ to emphasise the application of a threshold.
@@
    $\text{Ga}(\nu;\,a=3,b=0.1,\mu=1)$, a thresholded Gamma distribution with shape $a=3$. This distribution has mode at $20$ and puts more probability mass away from the interval $(1,10)$ as compared to the first three prior options. Thus, it promotes larger values of $\nu$ resulting in the smooth behavior rather than piecewise constant. This can be beneficial, if the task is to reconstruct smooth features in the object rather than sharp edges.
@@ 3.3 Posterior under the Student’s $t$ prior
The Bayesian inverse problem of estimating the posterior eq. 6b under the hierarchical prior in eq. 13a is written as the hierarchical Bayesian inverse problem of determining the posterior density
@@
In particular, the full conditional densities for each parameter are
@@
Now we can replace the corresponding densities on each conditional to derive them in closed form when possible.
Conditional 1. The density on $\bm{x}\;|\;\sigma_{\mathrm{obs}}^{2},\tau^{2},\bm{w}^{2}$ sets a linear Gaussian Bayesian inverse problem which has closed form solution for the posterior (see, e.g., (Kaipio and Somersalo, 2005, p.78)). Therefore
$p\left(\bm{x}\;|\;\sigma_{\mathrm{obs}}^{2},\tau^{2},\bm{w}^{2}\right)$ $\propto\exp\left(-\frac{1}{2}\left(\frac{1}{\sigma_{\mathrm{obs}}^{2}}\left\|\bm{y}-\bm{A}\bm{x}\right\|_{2}^{2}+\bigl{\|}\bm{\Lambda}^{\nicefrac{{1}}{{2}}}(\tau^{2},\bm{w}^{2})\,\bm{x}\bigr{\|}_{2}^{2}\right)\right)$ (18a)
$\bm{x}\;|\;\sigma_{\mathrm{obs}}^{2},\tau^{2},\bm{w}^{2}$ $\sim\mathcal{N}\left(\widetilde{\bm{\mu}}(\tau^{2},\bm{w}^{2})\,,\,\widetilde{\bm{\Lambda}}^{-1}(\tau^{2},\bm{w}^{2})\right),$ (18b)
@@
recall the prior precision matrix $\bm{\Lambda}(\tau^{2},\bm{w}^{2})$ is defined in eq. 12. The Cholesky factorization provides the most direct sampling algorithm for a Gaussian distribution. In this approach, a sample is obtained as $\bm{x}^{\star}=\widetilde{\bm{\mu}}+\widetilde{\bm{\Lambda}}^{-\nicefrac{{1}}{{2}}}\bm{u}$, where $\bm{u}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ represents a standard Gaussian random vector, and $\widetilde{\bm{\Lambda}}^{\nicefrac{{1}}{{2}}}$ denotes a lower triangular matrix with positive real diagonal entries. However, the Cholesky strategy becomes computationally prohibitive due to the need for factorizing the matrix $\widetilde{\bm{\Lambda}}(\tau^{2},\bm{w}^{2})$. This difficulty also arises in MCMC algorithms for Gaussian distributions, such as preconditioned Crank–Nicolson Cotter et al. (2013). To overcome this, we can use Krylov subspace methods to sample from the Gaussian conditional distribution in section 3.3. In particular, we can draw a sample from $p\left(\bm{x}\;|\;\sigma_{\mathrm{obs}}^{2},\tau^{2},\bm{w}^{2}\right)$ through randomized optimization. For further details on this formulation, refer to (Bardsley, 2019, p.81).
@@
Conditional 2. The conditional density for the noise variance $\sigma_{\mathrm{obs}}^{2}$ is obtained as the product of Gaussian and inverse-gamma densities. Through conjugate relations, this results is another inverse-gamma density that can be directly simulated
$p(\sigma_{\mathrm{obs}}^{2}\;|\;\bm{x})$ $\propto p(\bm{y}\;|\;\bm{x},\sigma_{\mathrm{obs}}^{2})p(\sigma_{\mathrm{obs}}^{2})$ (20a)
@@
$\propto(\sigma_{\mathrm{obs}}^{2})^{-\nicefrac{{m}}{{2}}}\exp\left(-\frac{1}{2\sigma_{\mathrm{obs}}^{2}}\left\|\bm{y}-\bm{A}\bm{x}\right\|_{2}^{2}\right)\left[(\sigma_{\mathrm{obs}}^{2})^{-a-1}\exp\left(-\frac{b}{\sigma_{\mathrm{obs}}^{2}}\right)\right]$ (20b)
$\propto(\sigma_{\mathrm{obs}}^{2})^{-(\nicefrac{{m}}{{2}}+a)-1}\exp\left(-\left[\frac{1}{2}\left\|\bm{y}-\bm{A}\bm{x}\right\|_{2}^{2}+b\right]\frac{1}{\sigma_{\mathrm{obs}}^{2}}\right)$ (20c)
$\sigma_{\mathrm{obs}}^{2}\;|\;\bm{x}$ $\sim\text{IG}\left(\frac{m}{2}+a,\frac{1}{2}\left\|\bm{y}-\bm{A}\bm{x}\right\|_{2}^{2}+b\right),$ (20d)
recall the hyperprior shape and scale parameters can be fixed as $a=1$ and $b=10^{-4}$, respectively.
Conditional 3. The conditional density for the squared-scale parameter $\tau^{2}$ is another inverse-gamma density
$p(\tau^{2}\;|\;\bm{x},\bm{w}^{2})$ $\propto p(\bm{x}\;|\;\tau^{2},\bm{w}^{2})p({\tau}^{2})$ (21a)
$\propto(\tau^{2})^{-\nicefrac{{k}}{{2}}}\exp\left(-\sum_{i=1}^{k}\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2\tau^{2}w_{i}^{2}}\right)\left[(\tau^{2})^{-a-1}\exp\left(-\frac{b}{\tau^{2}}\right)\right]$ (21b)
$\propto(\tau^{2})^{-(\nicefrac{{k}}{{2}}+a)-1}\exp\left(-\left[\sum_{i=1}^{k}\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2w_{i}^{2}}+b\right]\frac{1}{\tau^{2}}\right)$ (21c)
$\tau^{2}\;|\;\bm{x},\bm{w}^{2}$ $\sim\text{IG}\left(\frac{k}{2}+a,\sum_{i=1}^{d}\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2w_{i}^{2}}+b\right),$ (21d)
@@
where $k=\{d,2d\}$ in one- and two-dimensional problems, respectively. Recall the hyperprior shape and scale parameters are the same as those of the noise variance.
Conditional 4. Since the auxiliary parameters are independent, we can derive the conditional density for each component of $\bm{w}^{2}$ as follows
$p(w_{i}^{2}\;|\;\bm{x},\tau^{2},\nu)$ $\propto p(\bm{x}\;|\;\tau^{2},\bm{w}^{2})p(\bm{w}^{2}\;|\;\nu)$ (22a)
@@
$\propto(w_{i}^{2})^{-\nicefrac{{1}}{{2}}}\exp\left(-\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2\tau^{2}w_{i}^{2}}\right)\left[(w_{i}^{2})^{-\nicefrac{{\nu}}{{2}}-1}\exp\left(-\frac{\nu}{2w_{i}^{2}}\right)\right]$ (22b)
@@
$\propto(w_{i}^{2})^{-(\nicefrac{{1}}{{2}}+\nicefrac{{\nu}}{{2}})-1}\exp\left(-\left[\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2\tau^{2}}+\frac{\nu}{2}\right]\frac{1}{w_{i}^{2}}\right)$ (22c)
$w_{i}^{2}\;|\;\bm{x},\tau^{2},\nu$ $\sim\text{IG}\left(\frac{\nu+1}{2},\,\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2\tau^{2}}+\frac{\nu}{2}\right).$ (22d)
Conditional 5. Finally, the conditional density for the degrees of freedom $\nu$ is written as

### v2

@@ 2.3 Gaussian scale mixtures
is a GSM with mixing density $p_{\scriptscriptstyle Z}(z)$ and Gaussian component density $p_{\scriptscriptstyle X|Z}(x\;|\;z)$.
Necessary and sufficient conditions for the existence of a GSM representation are discussed in [10]. They essentially amount to some positivity condition on the derivatives of $p_{\scriptscriptstyle X}(x)$, and also $p_{\scriptscriptstyle X}(x)$ has to be both symmetric and unimodal.
Random variables written as GSMs are usually referred to as sub-Gaussian, i.e., their distributions have strong tail decay, at least as fast as the tails of a Gaussian [41, p.77]. For example, many important symmetric random variables can be constructed using a GSM: the logistic and Laplace distributions [10], generalized Gaussian distributions [42], as well as, the Student’s $t$ and $\alpha$-stable families of distributions [43, p.173]. We summarize some of these constructions in Table 1.
Distribution Gaussian Mixture Parameters
$p_{\scriptscriptstyle X}(x)$ $p_{\scriptscriptstyle X|Z}(x\;|\;z)$ $p_{\scriptscriptstyle Z}(z)$
Laplace: $\mathcal{L}\left(\mu,\tau\right)$ $\mathcal{N}(\mu,z)$ Exponential: $\mathcal{E}\left(\lambda\right)$ rate:
$\lambda=1/(2\tau^{2})$
Student’s $t$ : $\mathcal{T}_{\nu}\left(\mu,\tau\right)$ $\mathcal{N}(\mu,\tau^{2}z)$ Inverse-gamma: $\text{IG}\left(\alpha,\beta\right)$ shape and rate:
$\alpha=\beta=\nu/2$
$\alpha$-stable: $\mathcal{S}_{\alpha}\left(0,\mu,\tau\right)$ $\mathcal{N}(\mu,2\tau^{2}z)$ $\alpha$-stable: $\mathcal{S}_{\alpha/2}\left(\beta,\widetilde{\mu},\widetilde{\tau}\right)$ skewness, location, scale:
$\beta=1,\widetilde{\mu}=0,\widetilde{\tau}=1$
Table 1: Important examples of GSMs (see, e.g., [33]).
When GSMs are used as priors for Bayesian inference, they induce a hierarchical structure, which is conditionally Gaussian at the first layer, and the mixing random variable becomes a hyperparameter at the second layer (see, e.g., [33, 44]). This representation is ideal in iterative algorithms such as Markov chain Monte Carlo (MCMC), since upon the introduction of the auxiliary mixing variable, the target random variable becomes conditionally Gaussian. Standard and efficient techniques are available to deal with Gaussian distributions compared to those required to analyze the more involved sub-Gaussian distributions.
@@ 3 Student’s $t$ MRF prior based on Gaussian scale mixture representation
We introduce a prior that combines the MRF structure with the parametric family of Student’s $t$ distributions (see also [32, Ch.5]). This distribution is characterized by tail-heaviness, which is modulated through the degrees of freedom parameter. Subsequently, we outline a method to explore the posterior distribution under this prior, with a particular focus on the modeling of the degrees of freedom parameter.
@@ 3.1 Formulation of the prior
We investigate a parametric family of MRF priors designed to capture both sharp and smooth features. This family utilizes the Gosset distribution, commonly known as Student’s $t$ distribution or simply the $t$ distribution (we use these names interchangeably throughout the presentation). We employ this general family as a prior for each component in the difference random vector
@@
where $\mathcal{T}_{\nu}\left(\mu,\tau^{2}\right)$ denotes Student’s $t$ distribution, and it has density function [43, p.49]
@@
Notably, in the limit $\nu\to\infty$, the $t$ distribution approaches a Gaussian distribution. Conversely, when $\nu=1$, the distribution corresponds to a Cauchy distribution. It is worth noting that in practical applications, the degrees of freedom parameter is typically constrained to values greater than one or two, ensuring a well-defined mean or variance for the prior distribution, respectively. Consequently, to emphasize edge preservation, attention is directed towards distributions with lower degrees of freedom; in the opposite case, when aiming to represent smooth features, distributions with higher degrees of freedom become necessary. Intuitively, when using heavy-tailed densities we are increasing the probability of occurrence of large jump events (i.e., we are promoting large values of $\bm{u}$, thereby inducing sharpness). To illustrate the relationship between the degrees of freedom and the tail-heaviness of the $t$ distribution, we visually depict in Figure 1 the logarithm of the density function of the univariate standard $t$ distribution $\mathcal{T}_{\nu}\left(0,1\right)$ with increasing degrees of freedom.
Figure 1: Logarithm of the probability density of the univariate standard Student’s $t$ distribution $\mathcal{T}_{\nu}\left(0,1\right)$ for different values of the degrees of freedom parameter $\nu$. As $\nu\to\infty$, the $t$ distribution $\mathcal{T}_{\infty}\left(0,1\right)$ is equivalent to the univariate standard Gaussian distribution $\mathcal{N}(0,1)$.
Sampling from posterior distributions in high dimensions under the Student’s $t$ MRF prior is potentially cumbersome. To improve the efficiency of the prior implementation, one can utilize the GSM representation of Student’s $t$ distribution. As we anticipated in Table 1, the mixing distribution is in this case inverse-gamma [10]. By incorporating this property, we can transform the original prior model into a hierarchical one, where the target difference random vector follows (conditionally) a more tractable Gaussian distribution. Therefore, we can write the GSM prior density for the $i$-th component of difference random vector as follows
@@
where ${w}_{i}^{2}$ is the mixing random variable and $\text{IG}(a,b)$ denotes the inverse-gamma distribution with shape parameter $a$ and scale parameter $b$. The inverse-gamma probability density is defined as
$p(w_{i}^{2})=\frac{b^{a}}{\Gamma(a)}(w_{i}^{2})^{-a-1}\exp\left(-\frac{b}{w_{i}^{2}}\right);$ (11)
from the GSM formulation eq. 10b, we see that the shape and rate parameters are $a=b=\nu/2$.
By utilizing the GSM in Equation eq. 10b and the linear relation described in Equation eq. 5, we can establish a conditionally Gaussian MRF prior for the random vector $\bm{X}$. In Gaussian MRFs, the precision matrix plays a crucial role, as its sparsity structure determines the neighborhood system that explains conditional relationships among the elements of $\bm{X}$. In edge-preserving inverse problems, the precision matrix is constructed based on the differences between the elements of the parameter vector. Specifically, it depends on the matrix $\bm{L}$ in Equation eq. 5 (for detailed information, refer to [9, p.68]). Consequently, we can express the prior precision matrix as
@@
As a result, by using the GSM representation of the $t$ distribution on the random vector of differences $\bm{U}$, we establish a hierarchical prior for all the unknowns as follows
$p(\bm{x},\tau^{2},\bm{w}^{2},\nu)=p(\bm{x}\;|\;\tau^{2},\bm{w}^{2})p(\bm{w}^{2}\;|\;\nu)p(\nu)p(\tau^{2}),$ (13a)
@@
$\tau^{2}\sim\text{IG}\left(a,b\right),\quad{w}_{i}^{2}\;|\;\nu\sim\text{IG}\left(\nicefrac{{\nu}}{{2}},\nicefrac{{\nu}}{{2}}\right),\quad\nu\sim p(\nu),$ (13c)
@@
In Equation eq. 10b, each mixing variable $w_{i}^{2}$ appears to control the variance locally for each difference component $u_{i}$. Analogously, the scale-squared parameter $\tau^{2}$ controls the variance globally across the domain. Since this parameter is relevant in the prior specification, we assume it is uncertain and endow it with an inverse-gamma hyperprior $\tau^{2}\sim\text{IG}(a,b)$. Analogously to the discussion in Remark 2.1, we set the shape and scale parameters as $a=1$ and $b=10^{-4}$.
Remark 3.2.
The global-local scheme associated with the hierarchical prior eq. 13a shares similarities with the horseshoe prior [27], however, the hyperpriors for the local and global scale parameters are defined as half-Cauchy distributions for which the GSM representation has a different form. Moreover, the hierarchical prior eq. 13a is also related to the models discussed in [45, 46], which employ a generalized inverse-gamma distribution as hyperprior.
@@ 3.2 Modeling the degrees of freedom
Estimation of the degrees of freedom parameter $\nu$ of Student’s $t$ distribution has been actively studied [14, 15, 16]. The choice of prior distribution is known to be challenging as this parameter is typically poorly identified from the data [47].
We consider two popular prior choices [17, 20]:
@@
    $\text{Ga}(\nu;\,a=2,b=0.1,\mu=0)$, the gamma distribution with the shape parameter $a\in\mathbb{R}_{>0}$, rate parameter $b\in\mathbb{R}_{>0}$, and location parameter $\mu\in\mathbb{R}_{>0}$. The gamma probability density is defined as
@@
    $\text{Ga}(\nu;\,a=2,b=0.1,\mu=1)$, a thresholded Gamma distribution which is equivalent to the prior choice $1$, but now excluding values $\nu\leq 1$. We remark that in this case $\mu\neq 0$, and we further employ notation $\mathrm{Ga}_{\nu>\mu}(\nu;\,a,b)$ to emphasize the application of a threshold.
@@
    $\text{Ga}(\nu;\,a=3,b=0.1,\mu=1)$, a thresholded Gamma distribution with shape $a=3$. This distribution has a mode at $20$ and puts more probability mass away from the interval $(1,10)$ as compared to the first three prior options. Thus, it promotes larger values of $\nu$ resulting in smooth behavior rather than piecewise constant. This can be beneficial if the task is to reconstruct smooth features in the object rather than sharp edges.
@@ 3.3 Posterior under the Student’s $t$ prior
The Bayesian inverse problem of estimating the posterior eq. 6b under the hierarchical prior in Equation eq. 13a is written as the hierarchical Bayesian inverse problem of determining the posterior density
@@
In particular, using the relation between joint, conditional, and marginal distributions [43, p.157], one can verify that for each parameter, the full conditional densities associated with Equation eq. 16 are
@@
Now we can replace the corresponding densities on each conditional to derive them in a closed form when possible.
Conditional 1: $p\left(\bm{x}\;|\;\sigma_{\mathrm{obs}}^{2},\tau^{2},\bm{w}^{2}\right)$. This density can be seen as a linear Gaussian Bayesian inverse problem which has a closed-form solution for the posterior (see, e.g., [8, p.78]). Therefore, we obtain
$\begin{split}p\left(\bm{x}\;|\;\sigma_{\mathrm{obs}}^{2},\tau^{2},\bm{w}^{2}\right)&\propto\exp\left(-\frac{1}{2}\left(\frac{1}{\sigma_{\mathrm{obs}}^{2}}\left\|\bm{y}-\bm{A}\bm{x}\right\|_{2}^{2}+\bigl{\|}\bm{\Lambda}^{\nicefrac{{1}}{{2}}}(\tau^{2},\bm{w}^{2})\,\bm{x}\bigr{\|}_{2}^{2}\right)\right)\\ \bm{x}\;|\;\sigma_{\mathrm{obs}}^{2},\tau^{2},\bm{w}^{2}&\sim\mathcal{N}\left(\widetilde{\bm{\mu}}(\tau^{2},\bm{w}^{2})\,,\,\widetilde{\bm{\Lambda}}^{-1}(\tau^{2},\bm{w}^{2})\right),\end{split}$ (18)
@@
recall the prior precision matrix $\bm{\Lambda}(\tau^{2},\bm{w}^{2})$ is defined in Equation eq. 12. The Cholesky factorization provides the most direct sampling algorithm for a Gaussian distribution. In this approach, a sample is obtained as $\bm{x}^{\star}=\widetilde{\bm{\mu}}+\widetilde{\bm{\Lambda}}^{-\nicefrac{{1}}{{2}}}\bm{u}$, where $\bm{u}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ represents a standard Gaussian random vector, and $\widetilde{\bm{\Lambda}}^{\nicefrac{{1}}{{2}}}$ denotes a lower triangular matrix with positive real diagonal entries. However, the Cholesky strategy becomes computationally prohibitive due to the need for factorizing the matrix $\widetilde{\bm{\Lambda}}(\tau^{2},\bm{w}^{2})$. This difficulty also arises in MCMC algorithms for Gaussian distributions, such as preconditioned Crank–Nicolson [48]. To overcome this, we can use Krylov subspace methods to sample from the Gaussian conditional distribution eq. 18. In particular, we can draw a sample from $p\left(\bm{x}\;|\;\sigma_{\mathrm{obs}}^{2},\tau^{2},\bm{w}^{2}\right)$ through randomized optimization. For further details on this formulation, refer to [9, p.81].
@@
Conditional 2: $p(\sigma_{\mathrm{obs}}^{2}\;|\;\bm{x})$. The conditional density for the noise variance is proportional to the product of Gaussian and inverse-gamma densities. Through conjugate relations, such density is another inverse-gamma density that can be directly simulated
@@
$\begin{split}p(\sigma_{\mathrm{obs}}^{2}\;|\;\bm{x})&\propto p(\bm{y}\;|\;\bm{x},\sigma_{\mathrm{obs}}^{2})p(\sigma_{\mathrm{obs}}^{2})\\ &\propto(\sigma_{\mathrm{obs}}^{2})^{-\nicefrac{{m}}{{2}}}\exp\left(-\frac{1}{2\sigma_{\mathrm{obs}}^{2}}\left\|\bm{y}-\bm{A}\bm{x}\right\|_{2}^{2}\right)\left[(\sigma_{\mathrm{obs}}^{2})^{-a-1}\exp\left(-\frac{b}{\sigma_{\mathrm{obs}}^{2}}\right)\right]\\ &\propto(\sigma_{\mathrm{obs}}^{2})^{-(\nicefrac{{m}}{{2}}+a)-1}\exp\left(-\left[\frac{1}{2}\left\|\bm{y}-\bm{A}\bm{x}\right\|_{2}^{2}+b\right]\frac{1}{\sigma_{\mathrm{obs}}^{2}}\right)\\ \sigma_{\mathrm{obs}}^{2}\;|\;\bm{x}&\sim\text{IG}\left(\frac{m}{2}+a,\frac{1}{2}\left\|\bm{y}-\bm{A}\bm{x}\right\|_{2}^{2}+b\right),\end{split}$ (20)
@@
recall from Remark 2.1 that the shape and scale parameters of the noise variance hyperprior are defined as $a=1$ and $b=10^{-4}$, respectively.
Conditional 3: $p(\tau^{2}\;|\;\bm{x},\bm{w}^{2})$. The conditional density for the scale-squared parameter $\tau^{2}$ is another inverse-gamma density
@@
$\begin{split}p(\tau^{2}\;|\;\bm{x},\bm{w}^{2})&\propto p(\bm{x}\;|\;\tau^{2},\bm{w}^{2})p({\tau}^{2})\\ &\propto(\tau^{2})^{-\nicefrac{{k}}{{2}}}\exp\left(-\sum_{i=1}^{k}\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2\tau^{2}w_{i}^{2}}\right)\left[(\tau^{2})^{-a-1}\exp\left(-\frac{b}{\tau^{2}}\right)\right]\\ &\propto(\tau^{2})^{-(\nicefrac{{k}}{{2}}+a)-1}\exp\left(-\left[\sum_{i=1}^{k}\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2w_{i}^{2}}+b\right]\frac{1}{\tau^{2}}\right)\\ \tau^{2}\;|\;\bm{x},\bm{w}^{2}&\sim\text{IG}\left(\frac{k}{2}+a,\sum_{i=1}^{d}\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2w_{i}^{2}}+b\right),\end{split}$ (21)
where $k=\{d,2d\}$ in one- and two-dimensional problems, respectively. Recall that the hyperprior shape and scale parameters are the same as those of the noise variance (cf. Remark 3.1).
Conditional 4: $p(w_{i}^{2}\;|\;\bm{x},\tau^{2},\nu)$. Since the auxiliary parameters are independent, we can derive the conditional density for each component of $\bm{w}^{2}$ as follows
@@
$\begin{split}p(w_{i}^{2}\;|\;\bm{x},\tau^{2},\nu)&\propto p(\bm{x}\;|\;\tau^{2},\bm{w}^{2})p(\bm{w}^{2}\;|\;\nu)\\ &\propto(w_{i}^{2})^{-\nicefrac{{1}}{{2}}}\exp\left(-\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2\tau^{2}w_{i}^{2}}\right)\left[(w_{i}^{2})^{-\nicefrac{{\nu}}{{2}}-1}\exp\left(-\frac{\nu}{2w_{i}^{2}}\right)\right]\\ &\propto(w_{i}^{2})^{-(\nicefrac{{1}}{{2}}+\nicefrac{{\nu}}{{2}})-1}\exp\left(-\left[\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2\tau^{2}}+\frac{\nu}{2}\right]\frac{1}{w_{i}^{2}}\right)\\ w_{i}^{2}\;|\;\bm{x},\tau^{2},\nu&\sim\text{IG}\left(\frac{\nu+1}{2},\,\frac{([\bm{L}\bm{x}]_{i}-\mu)^{2}}{2\tau^{2}}+\frac{\nu}{2}\right).\end{split}$ (22)
Conditional 5: $p(\nu\;|\;\bm{w}^{2})$. Finally, the conditional density for the degrees of freedom parameter $\nu$ is written as